DYNAMICS OF LOCALIZED ANGULAR MOMENTUM AND MULTI-SURFACE ROTATIONAL ENERGY ANISOTROPY FOR INTERNAL-ROTOR MOLECULES AND POSSIBLE SYMMETRY CONVERSION EF-FECTS

\underline{JUSTIN C. MITCHELL}, WILLIAM G. HARTER, Department of Physics, University of Arkansas, Fayetteville, AR 72701.

Rigid or semi-rigid rotors with attached gyroscopes approximately model dynamics and spectra of molecules with internal rotation ^{a} or rovibrational coupling ^{b}. Classical dynamics of axially constrained rotors are approximated by intersecting rotational-energy-surfaces (RES) that have (J-S).B.(J-S) forms in the limit of constraints that do no work. Improved semiclassical eigen-RES have avoided crossings that make concentric nested surfaces whose anisotropy determines approximate levels, symmetries and dynamics ^{c}. Semi-classical eigensolutions are compared to those found by direct diagonalization. Possible application to spectra of methyl-rotor molecules are discussed.

aJ. Ortigoso, I. Kleiner, and J. Hougen, J. Chem. Phys. 110, 11688 (1999).

bH. S. Yoo, M. J. DeWitt, and B. H. Pate, J. Chem. Phys, 108, 1348 (2004).

cW. G. Harter, Comp. Phys. Reports 8, 320-393(1988) Sections 6.2-3.